# Privacy vs. Profit: The Impact of Google’s Manifest Version 3 (MV3) Update on Ad Blocker Effectiveness

### v1

@@ B.3 Effectiveness of Early MV3 Ad Blockers
We replicated the browser-based experiment using the first available versions of AdGuard MV3 (version 0.3.9) and uBlock MV3 (version 0.1.22.806) ad blockers. We excluded Adblock Plus and Stands ad blockers due to the availability of only a single version (Adblock Plus 4.5.1; Stands 2.1.10) in MV3, which we had already included in Table 1.
This robustness test thus compares the effectiveness of these early versions against their MV2 counterparts. Figure 9 presents independent t-test comparisons for the number of blocked ads (Panel A) and blocked trackers (Panel B).
Figure 9: Number of blocked ads and trackers between individual ad blockers and the MV3+ ad blocker group for early MV3 ad blocker versions compared to their MV2 counterparts. Panels A and C show independent t-test comparisons between the mean values of the early MV3 ad blockers (AdGuard MV3 version 0.3.9; uBlock MV3 version 0.1.22.806) and their MV2 instances (AdGuard MV2 version 4.2.168; uBlock MV2 version 1.52.0) for the number of blocked ads and trackers, respectively. We exclude the MV3+ ad blocker group from these comparisons. Multiplying the 4 browser instances (AdGuard MV3 version 0.3.9, uBlock MV3 version 0.1.22.806, AdGuard MV2 version 4.2.168; uBlock MV2 version 1.52.0) with the number of websites (114) yields the number of observations (N = 456). Panels B and D show comparisons between the MV3+ ad blocker group and the individual early MV3 ad blockers for the number of blocked ads and trackers, respectively. Multiplying the 3 browser instances (AdGuard MV3 version 0.3.9, uBlock MV3 version 0.1.22.806, and MV3+ ad blocker group) with the number of websites (114) yields the number of observations (N = 342). Error bars represent ±$\pm 1$ SEs.
For the number of blocked ads, we found no significant differences between the early MV3 and MV2 versions for both AdGuard and uBlock $\Delta=0.00$, $p=1.00$. Similarly, for the number of blocked trackers, early MV3 ad blockers did not differ significantly from their MV2 counterparts (for AdGuard, $\Delta=-0.12$, $p=0.96$; for uBlock, $\Delta=0.09$, $p=0.97$). Moreover, comparisons between the MV3+ ad blocker group and the early MV3 configurations revealed no significant differences in either measure.
@@ B.4 Effectiveness of MV3 Ad Blockers on Firefox
On Firefox, we found only uBlock available in both MV2 and MV3 instances. Because Firefox does not support the Azerion Ad Expert extension, we manually counted the number of blocked ads, and we used a custom extension to record the tracker data, mirroring our methodology for Chrome (see Section 5.2).
Figure 10 presents independent t-test comparisons for uBlock MV3 vs. MV2 instances on Firefox. For the number of blocked ads, the MV3 and MV3 instances showed no significant difference ($\Delta=-0.01$, $p=0.97$). For the number of blocked trackers, we also found no significant differences between the MV3 and MV2 instances ($\Delta=-0.24$, $p=0.92$).
Figure 10: Number of blocked ads and trackers between uBlock MV3 and MV2 on Firefox. Panel A shows independent t-test comparison between the mean values of the MV3 and MV2 instances of uBlock ad blocker for the number of blocked ads on Firefox. We exclude the MV3+ ad blocker group from this comparison. Multiplying the 2 browser instances (uBlock MV2, uBlock MV3) with the number of websites (114) yields the number of observations (N = 228). Panel B shows independent t-test comparison between the mean values of the MV3 and MV2 instances of uBlock ad blocker for the number of blocked trackers on Firefox. We also exclude the MV3+ ad blocker group from this comparison. Multiplying the 2 browser instances (uBlock MV2, uBlock MV3) with the number of websites (114) yields the number of observations (N = 228). Error bars represent ±$\pm 1$ SEs.

### v2

@@ B.3 Effectiveness of Early MV3 Ad Blockers
We replicated the browser-based experiment using our main sample of 1,000 websites on July 20, 2025 using the first available versions of AdGuard MV3 (version 0.3.9) and uBlock MV3 (version 0.1.22.806) ad blockers. We excluded Adblock Plus and Stands ad blockers due to the availability of only a single version (Adblock Plus 4.5.1; Stands 2.1.10) in MV3, which we had already included in Table 1.
This robustness test thus compares the effectiveness of these early versions against their MV2 counterparts. Figure 9 presents independent t-test comparisons for the number of blocked ads (Panels A and B) and blocked trackers (Panels C and D).
Figure 9: Number of blocked ads and trackers between individual ad blockers and the MV3+ ad blocker group for early MV3 ad blocker versions compared to their MV2 counterparts. Panels A and C show independent t-test comparisons between the mean values of the early MV3 ad blockers (AdGuard MV3 version 0.3.9; uBlock MV3 version 0.1.22.806) and their MV2 instances (AdGuard MV2 version 4.2.168; uBlock MV2 version 1.52.0) for the number of blocked ads and trackers, respectively. We exclude the MV3+ ad blocker group from these comparisons. Multiplying the 4 browser instances (AdGuard MV3 version 0.3.9, uBlock MV3 version 0.1.22.806, AdGuard MV2 version 4.2.168; uBlock MV2 version 1.52.0) with the number of websites (939) yields the number of observations (N = 3,756). Panels B and D show comparisons between the MV3+ ad blocker group and the individual early MV3 ad blockers for the number of blocked ads and trackers, respectively. Multiplying the 3 browser instances (AdGuard MV3 version 0.3.9, uBlock MV3 version 0.1.22.806, and MV3+ ad blocker group) with the number of websites (939) yields the number of observations (N = 2,817). Values shown are from a single measurement run (no averaging); error bars represent $\pm 1$ SE across websites.
For the number of blocked ads, we found no significant differences between the early MV3 and MV2 versions for both AdGuard and uBlock. Similarly, for the number of blocked trackers, early MV3 ad blockers did not differ significantly from their MV2 counterparts. Moreover, in comparisons between the MV3+ ad blocker group and the early MV3 ad blockers, we observe no differences for blocked ads for both AdGuard and uBlock. For blocked trackers, the MV3+ ad blocker group blocks more than AdGuard MV3 ($\Delta=4.43,p<0.001$), while uBlock shows no difference.
@@ B.4 Effectiveness of MV3 Ad Blockers on Firefox
On Firefox, we found only uBlock available in both MV3 and MV2 instances. Because Firefox does not support the Azerion Ad Expert extension, we could not count the number of blocked ads according to our methodology for Chrome (see Section 5.2). For the number of blocked trackers, however, we used a custom extension to record the tracker data, mirroring our methodology for Chrome.
We initiated the Firefox browser-based experiment on a main sample of 1,000 websites across five separate measurement runs on July 12, 2025, out of which we consistently obtained results for 824 websites. We then averaged those results across five runs for our measures of interest: the HTML file size (kb) of loaded websites and the number of blocked trackers.
Table 6 reports descriptive and t-test statistics for the HTML file size (kb) of loaded websites under the usage of uBlock MV3 and MV2 on Firefox. The results of this test show no significant differences between the MV3 and MV2 instances of uBlock ad blocker in HTML file size.
Table 6: HTML file size (kb) comparison for ad-blocking and anti-tracking effectiveness between uBlock MV3 and MV2 on Firefox. Multiplying the 2 browser instances (uBlock MV2, uBlock MV3) with the number of websites (824) yields the number of observations (N = 1,648). Values shown are per-website averages across five runs.
Condition MV2 MV3 Comparison
N (each) mean SD mean SD t-stat $p$-val
uBlock 824 413.47 428.77 409.66 421.24 0.18 0.86
Figure 10 presents independent t-test comparisons for uBlock MV3 vs. MV2 instances on Firefox. For the number of blocked trackers, we found no significant differences between the MV3 and MV2 instances ($\Delta=-0.05$, $p=0.93$).
Figure 10: Number of trackers between uBlock MV3 and MV2 on Firefox. This figure shows independent t-test comparison between the mean values of the MV3 and MV2 instances of uBlock ad blocker for the number of blocked trackers on Firefox. We exclude the MV3+ ad blocker group from this comparison. Multiplying the 2 browser instances (uBlock MV2, uBlock MV3) with the number of websites (824) yields the number of observations (N = 1,648). Values shown are per-website averages across five runs; error bars represent $\pm 1$ SE across websites.
@@
B.5 Visual Inspection of MV3 vs. MV2 Ad Blocker Screenshots
We conducted a visual inspection to compare ad-blocking outcomes under MV3 versus MV2. From our main sample of 924 websites, we drew a random subset of 100 website visits. For each selected site, we placed the MV3 and the corresponding MV2 of the same ad blocker side by side and compared the resulting screenshots. For example, we might have randomly picked cnn.com and, conditional on that draw, randomly chose uBlock; we then compared the cnn.com screenshots produced by uBlock MV3 and uBlock MV2. The premise is that if MV3 introduces issues that are not present under MV2, systematic visual differences will emerge between the two screenshots.
As discussed in Section 3.1, we used the following definitions in the visual assessment:
Ad flickering. A brief, visible appearance of ads on a page before they are hidden or removed by the ad blocker. In screenshots, ad flickering should manifest as discrepancies between visible ad counts in the images and our automated ad counts. If an ad flickers but the automation misses it, the flickering ad should sometimes still be visible in the screenshots.
Breakage. Unintended blocking of legitimate (non-ad) content by an ad blocker. Breakage is recorded when one manifest version blocks legitimate content that remains visible under the other manifest version.
Cosmetic placeholders. Residual empty spaces or visual artifacts left after ad removal (e.g., blank containers), indicating incomplete cosmetic filtering. These are recorded when such empty white spaces are visible under one manifest version but not the other. The screenshots in Figure 11 showcase one example of our comparison that reveals such a leftover cosmetic placeholder.
(a) Screenshot using the uBlock Origin MV2 ad blocker
(b) Screenshot using the uBlock Origin Lite MV3 ad blocker
Figure 11: Screenshots showcasing a leftover cosmetic placeholder covering an ad under MV3, that was previously not visible under MV2, on the website chocolatecoveredkatie.com using uBlock Origin MV2 and uBlock Origin Lite MV3.
Two independent human evaluators assessed each outcome based on a specific question (one question per outcome), and compared the MV3 and MV2 screenshots:
  • “Do you see any discrepancy in visible ads in the provided screenshots versus the provided ad counts?” (ad flickering)
  • “Do you see any legitimate (non-ad) content being blocked under one manifest version that is not blocked under the other?” (breakage)
  • “Do you see any empty white spaces visible under one manifest version that are not visible under the other?” (cosmetic placeholders)
Evaluators reviewed the MV3 and MV2 screenshots side by side and answered the three questions above. We report results as Evaluator 1 (Evaluator 2). The two evaluators found no substantial ad-flickering discrepancies between MV3 and MV2: 4 (4) missed ads, split evenly across MV3 and MV2. They also observed no major breakage differences: 5 (5) MV3-only versus 5 (3) MV2-only cases. However, they found that MV3 left over cosmetic placeholders in 21% (22%) of cases versus 0% (0%) under MV2.
Within this 100-site visual audit, we did not observe meaningful MV3-specific ad flickering or systematic breakage relative to MV2. However, incomplete cosmetic filtering were notably more prevalent under MV3, with cosmetic placeholders appearing in roughly one-fifth of comparisons under MV3 and in none under MV2. Although our visual inspection is limited by the static nature of our screenshots, these findings suggest that, at present, MV3’s most visible shortcoming relative to MV2 is cosmetic rather than functional ad removal.